OpenStudy (anonymous):

Which expression is equivalent to 1 – cos^2 x?

OpenStudy (anonymous):

\[\bf \sin^2(x)+\cos^2(x)=1\]

OpenStudy (anonymous):

Can you figure it out from here?

OpenStudy (mayankdevnani):

always remember this identity: \[\huge \sin^2x+\cos^2x=1\]

OpenStudy (mayankdevnani):

so, \[\huge \cos^2x=???\]

OpenStudy (mayankdevnani):

can you solve it?? @thardyyyy

OpenStudy (anonymous):

Re-arrange the identity.

OpenStudy (mayankdevnani):

@thardyyyy are you there?